OpenStudy (anonymous):

Catapults are being used to hit an island. At 20 kilometers from shore the incoming boulders were located at 2800 meters of altitude. At 40 kilometers from shore, they were spotted at 4200 meters of altitude. And at 70 kilometers out, the boulders were passing at 5550 meters of altitude. How far from their shore is their enemies catapult? Are their shots too short or too long to hit us and by how much?

OpenStudy (anonymous):

We need to use the equation of a parabola here? do you know it?

OpenStudy (anonymous):

y = ax² + bx + c

OpenStudy (anonymous):

I believe thats it

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now lets use that to solve the problem

OpenStudy (anonymous):

let x and y be in kilimoters

OpenStudy (anonymous):

we have three points to work with

OpenStudy (anonymous):

(20,2.8) (40,4.2) (70,5.5)

OpenStudy (anonymous):

We must use these point to solve: y = ax² + bx + c

OpenStudy (anonymous):

are you paying attention or not?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now i will set up the equations, and you must solve them or at least try, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c + 20b + 400a = 2.8, c + 40b + 1600a = 4.2, and c + 70b + 4600a = 5.5.

OpenStudy (anonymous):

which am I solving for a, b, or c

OpenStudy (anonymous):

now you understand how i got these right, i places the x coordinated from our 3 points ((20,2.8) (40,4.2) (70,5.5)) that means the 20 the 40 and 70 replaced the x's in the equation y = ax² + bx + c

OpenStudy (anonymous):

well, we are going to solve these like we would solve a system of equations

OpenStudy (anonymous):

yes I get stuck right here where we are now

OpenStudy (anonymous):

now subtract the 1st equation from the second please

OpenStudy (anonymous):

and then tell me what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

20b+1200a=1.4

OpenStudy (anonymous):

okay, good now subtract the 1st equation from the third, and tell me what you get

OpenStudy (anonymous):

50b+4200a=2.7

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now subtract 5/2 of the 1st equation from the second

OpenStudy (anonymous):

I dont get it what do u mean subtract 5/2

OpenStudy (anonymous):

Ok I subtract 5/2 from 1200a correct but when I do it. It doesn't look right

OpenStudy (anonymous):

Multiply 1200, by 2.5 and subtract it from 4200

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The multiply 20 by 2.5 and subtract it from 50

OpenStudy (anonymous):

Then multiply 1.4 by 2.5 and subtract it from 2.7

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

I got -2950

OpenStudy (anonymous):

you should have gotten 1200a=-0.8

OpenStudy (anonymous):

Check this out

OpenStudy (anonymous):

ok hold on trying it again

OpenStudy (anonymous):

ok 1200*2.5=3000 4200-3000=1200

OpenStudy (anonymous):

(2.5)(20b+1200a=1.4)=50b+3000a=3.5

OpenStudy (anonymous):

now subtract that from our second equation: 50b+4200a=2.7

OpenStudy (anonymous):

we get: 50b+4200a=2.7 - 50b+3000a=3.5 ---------------- =

OpenStudy (anonymous):

1200a=-0.8

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now lets solve for a

OpenStudy (anonymous):

how can we do that?

OpenStudy (anonymous):

thats right, we divide both sides of 1200a=-0.8, by 1200 and get?

OpenStudy (anonymous):

a=-0.00066667

OpenStudy (anonymous):

oh ok if we divided both sides by 1200 we get a=-0.8?

OpenStudy (anonymous):

i mean a=-8/12,000

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

not* you see how i changed -0.8 to -8 and 1200 to 12000

OpenStudy (anonymous):

note*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, now we have found a, lets find c and b. To find b, simply replace a with -8/12000 in the equation 50b + 4200a = 2.7

OpenStudy (anonymous):

and solve, tell me what you get

OpenStudy (anonymous):

0.11

OpenStudy (anonymous):

b=0.11

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now to find c, simple replace a with -8/12000 and b with 0.11(or 5.5/50) in the equation c + 20b + 400a = 2.8

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

c=13/15

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I did as u said c+20(0.11)+400(-8/12000)=2.8

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now we have all three values for a,b and c

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now to determine where the boulders will hit solve ax² + bx + c = 0.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This will give give the x values for where it was launched from and where it will hit. Using these values and the fact that the place you are at is at (0,0), the distance from you can be found.

OpenStudy (anonymous):

now: If the x value is positive, the boulders fell short. If the x value is negative, the boulder went too far

OpenStudy (anonymous):

x=172.5347,-7.534716

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

very good

OpenStudy (anonymous):

Thank you very very much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Are u still here I have one more question

OpenStudy (anonymous):

i hope its not as long as this one :)

OpenStudy (anonymous):

lol no but if the questions ask me How far from their shore is their enemies catapult? Are their shots too short or too long to hit us and by how much? How would I plug in my answers here?

OpenStudy (anonymous):

I hate word problems they confuse me

OpenStudy (anonymous):

right we answered this already right?

OpenStudy (anonymous):

Yes just not sure how to respond if x=172.5347,-7.534716

OpenStudy (anonymous):

if x is positive there shots fell short, if x is negative there shots went to far

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok got it thanks again